# Learning Where to Learn: Training Distribution Selection for Provable OOD Performance

### v1

@@ E.3 NtD linear operator learning
$x\mapsto a(x)=\exp\Biggl{(}\sum_{j=1}^{M+1}\sum_{k=1}^{M+1}\frac{2\sigma}{\bigl{(}(j\pi)^{2}+(k\pi)^{2}+\tau^{2}\bigr{)}^{\alpha/2}}z_{jk}\sin(j\pi x_{1})\sin(k\pi x_{2})\Biggr{)}$ (E.11)
@@
with fixed parameters $\alpha=2$, $\tau=3$, $\sigma=3$ and random coefficients $z_{jk}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. Once these random coefficients are drawn, they are fixed throughout the example. Thus, in this experiment, our realization of the conductivity $a$ remains fixed; see the left panel of Fig. 3 for a visualization of this $a$. Our goal is to develop neural operator approximations of the NtD linear operator $\mathcal{G}^{\star}\coloneqq\Lambda_{a}^{-1}$ that maps the Neumann boundary condition to the Dirichlet boundary condition according to the PDE (D.2) (with Neumann data $g$ given instead of Dirichlet data $f$). Using the alternating minimization algorithm (AMA, Alg. 2) from Sec. 4.2, we want to optimize the training distribution over the Neumann boundary conditions to achieve small OOD generalization error (3.3).
@@
To evaluate the OOD generalization error, we need to assign a random measure $\mathbb{Q}$. Let $\mathbb{Q}=\frac{1}{K}\sum_{k=1}^{K}\delta_{\nu^{\prime}_{k}}$, where $\{\nu_{k}^{\prime}\}_{k=1}^{K}$ are distributions over functions. For a fixed $k$, the function $g\sim\nu_{k}^{\prime}$ is a Neumann boundary condition. It takes the form $g\colon\partial\Omega\to\mathbb{R}$, where
@@
$t\mapsto g\bigl{(}r(t)\bigr{)}=10\sin\mathopen{}\mathclose{{}\left(\omega_{k}\mathopen{}\mathclose{{}\left(t-\tfrac{1}{2}}\right)}\right)+\sum_{j=1}^{N_{\mathrm{basis}}}\frac{\sigma_{k}}{\bigl{(}(j\pi)^{2}+\tau_{k}^{2}\bigr{)}^{\alpha_{k}/2}}z_{j}\sqrt{2}\sin(j\pi t)\,.$ (E.12)
@@
We view $r\colon[0,1]\to\partial\Omega$ as a 1D parametrization of the boundary $\partial\Omega\subset\mathbb{R}^{2}$ for the domain $\Omega\coloneqq(0,1)^{2}$. The entire domain is discretized on an $M\times M$ grid, where $M$ is as in (E.11). Here $z_{j}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$, $\alpha_{k}=1.5$ for all $k$, and $N_{\mathrm{basis}}=4M-4$. We set the grid parameter to be $M=128$ and sample $K=64$ test distributions in this example. The parameters $\omega_{k}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\rho([-2\pi,2\pi])$ and $\tau_{k}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\rho([0,50])$ are sampled from a custom distribution $\rho=\rho([x_{\min},x_{\max}])$ and held fixed throughout the algorithm. The custom distribution $\rho$ has the probability density function
$p(x)=C\sin\bigl{(}x-\tfrac{\pi}{2}\bigr{)}+1$ (E.13)
@@
for $x\in[x_{\min},x_{\max}]\subset\mathbb{R}$, where $C$ is a normalizing constant ensuring that $p$ integrates to $1$ over the domain $[x_{\min},x_{\max}]$. Samples are generated by using inverse transform sampling. Any time a Neumann boundary condition $g$ is sampled according to (E.12) and (E.13), we subtract the spatial mean of $g$ from itself to ensure that the resulting function integrates to zero. This is necessary to make the Neumann PDE problem—as well as the NtD operator itself—well-posed.
@@
$t\mapsto\widetilde{g}(t)=10\sin\mathopen{}\mathclose{{}\left(\omega\mathopen{}\mathclose{{}\left(t-\tfrac{1}{2}}\right)}\right)+\sum_{j=1}^{N_{\mathrm{basis}}}\frac{\sigma}{\bigl{(}(j\pi)^{2}+\tau^{2}\bigr{)}^{\alpha/2}}\widetilde{z}_{j}\sqrt{2}\sin(j\pi t)\,,$ (E.14)
where $\widetilde{z}_{j}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$, $\alpha=2$, $\tau=3$, and $\sigma=\tau^{3/2}$. That is, $\nu=\operatorname{Law}(\widetilde{g})$ with $\widetilde{g}$ as in (E.14). The frequency parameter $\omega$ appearing in the mean term of (E.14) is optimized using Alg. 2 to identify a training distribution that minimizes average OOD generalization error with respect to $\mathbb{Q}$.
Note that $g$ and $\widetilde{g}$ are finite-term truncations of the KLE of Gaussian measures with Matérn-like covariance operators; recall Lem. A.2. In particular, for $\widetilde{g}$, and similarly for $g$, the covariance operator of the Gaussian measure has eigenpairs
$\lambda_{j}=\frac{\sigma^{2}}{\bigl{(}(j\pi)^{2}+\tau^{2}\bigr{)}^{\alpha}}\quad\text{and}\quad t\mapsto\phi_{j}(t)=\sqrt{2}\,\sin(j\pi t)\quad\text{for}\quad j=1,2,\dots,N_{\mathrm{basis}}\,.$ (E.15)
Using (E.15) and (A.1), the squared $2$-Wasserstein distance in Prop. 3.1 reduces to
@@
[…]
These closed-form reductions follow directly from the finite-dimensional KLE and Ex. A.1.
@@
$\max_{(g_{1},g_{2})\in\mathsf{S}}\frac{\lVert\mathcal{G}(g_{1})-\mathcal{G}(g_{2})\rVert_{L^{2}(\partial\Omega)}}{\lVert g_{1}-g_{2}\rVert_{L^{2}(\partial\Omega)}}$ (E.18)
@@
The model architecture is a DeepONet comprising a branch network with five fully connected layers of sizes $44M-$ 4, $512$, $512$, $512$, and $200$, and a trunk network with five fully connected layers of sizes $2$, $128$, $128$, $128$, and $200$, respectively; see SM D.1 for details. All layers use ReLU activation functions. Weights are initialized using the Glorot normal initializer. The model is trained with the Adam optimizer with a learning rate of $10^{-3}$. After the model-training half step in the alternating minimization scheme Alg. 2, the parameter $\omega$ in (E.14) is optimized using SciPy’s default implementation of differential evolution with a tolerance of $10^{-7}$. Due to the presence of many local minima in the optimization landscape, this global optimization algorithm is necessary.
@@
The initial value for $\omega$ is set to $8\pi$. For every model-training step, $500$ pairs of Neumann and Dirichlet data are generated using the current value of $\omega$. For practical reasons, the model is initially trained for $2000$ iterations of Adam. If the AMA loss has not improved relative to the previous iteration, training continues for an additional $2000$ iterations. This process is repeated up to $10$ times, after which the algorithm proceeds to the parameter optimization step for the training distribution $\nu$. During distribution parameter optimization, the AMA loss is evaluated using $500$ newly generated data pairs, used specifically to compute the first term of the objective (4.7). To ensure deterministic behavior, the $500$ sets of random coefficients $\{z_{j}\}_{j=1}^{N_{\mathrm{basis}}}$ are fixed separately for training and optimization. This results in $1000$ total sets of coefficients fixed for the entire algorithm. This process is repeated across $80$ independent runs using the same conductivity but different realizations of the random coefficients $\{z_{j}\}$. The relative AMA loss and relative OOD error is computed for each run; see (E.19).
@@
Relative OOD $\text{Relative OOD Error}\coloneqq\frac{\operatorname{\mathbb{E}}_{\nu^{\prime}\sim\mathbb{Q}}\operatorname{\mathbb{E}}_{g\sim\nu^{\prime}}\|\mathcal{G}^{\star}g-\mathcal{G}_{\theta}(g)\|_{L^{2}(\partial\Omega)}^{2}}{\operatorname{\mathbb{E}}_{\nu^{\prime}\sim\mathbb{Q}}\operatorname{\mathbb{E}}_{g\sim\nu^{\prime}}\|\mathcal{G}^{\star}g\|_{L^{2}(\partial\Omega)}^{2}}\,,$ (E.19)
@@
Relative OOD $\text{Relative OOD Error}\approx\frac{\frac{1}{K}\sum_{k=1}^{K}\frac{1}{N}\sum_{n=1}^{N^{\prime}}\|\mathcal{G}^{\star}g_{k,n}-\mathcal{G}_{\theta}(g_{k,n})\|_{L^{2}(\partial\Omega)}^{2}}{\frac{1}{K}\sum_{k=1}^{K}\frac{1}{N^{\prime}}\sum_{n=1}^{N^{\prime}}\|\mathcal{G}^{\star}g_{k,n}\|_{L^{2}(\partial\Omega)}^{2}}\,,$ (E.20)
@@
The computational resources here are the same as those reported in SM E.2, except three total CPU and GPU hours were needed for all independent runs.
E.4 Darcy flow forward operator learning
@@
In this final example from Sec. 5, we are interested in learning another operator that is relevant and related to the EIT problem. In contrast with the previous example in SM E.3 where we fix the conductivity, the Darcy flow forward operator that we consider is nonlinear and maps the conductivity function $a$ to the PDE solution $u$ of

### v3

@@ F.3 NtD linear operator learning
$x\mapsto a(x)=\exp\Biggl(\sum_{j=1}^{M+1}\sum_{k=1}^{M+1}\frac{2\sigma}{\bigl((j\pi)^{2}+(k\pi)^{2}+\tau^{2}\bigr)^{\alpha/2}}z_{jk}\sin(j\pi x_{1})\sin(k\pi x_{2})\Biggr)$ (F.11)
@@
with fixed parameters $\alpha=2$, $\tau=3$, $\sigma=3$ and random coefficients $z_{jk}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. Once these random coefficients are drawn, they are fixed throughout the example. Thus, in this experiment, our realization of the conductivity $a$ remains fixed; see the left panel of Fig. 3 for a visualization of this $a$. Our goal is to develop neural operator approximations of the NtD linear operator $\mathcal{G}^{\star}\coloneqq\Lambda_{a}^{-1}$ that maps the Neumann boundary condition to the Dirichlet boundary condition according to the PDE (F.6) (with Neumann data $g$ given instead of Dirichlet data $f$). Using the alternating minimization algorithm (AMA, Alg. 2) from Sec. 4.2, we want to optimize the training distribution over the Neumann boundary conditions to achieve small OOD error (3.3).
@@
To evaluate the OOD error, we need to assign a random measure $\mathbb{Q}$. Let $\mathbb{Q}=\frac{1}{K}\sum_{k=1}^{K}\delta_{\nu^{\prime}_{k}}$, where $\{\nu_{k}^{\prime}\}_{k=1}^{K}$ are distributions over functions. For a fixed $k$, the function $g\sim\nu_{k}^{\prime}$ is a Neumann boundary condition. It takes the form $g\colon\partial\Omega\to\mathbb{R}$, where
@@
$t\mapsto g\bigl(r(t)\bigr)=10\sin\mathopen{}\mathclose{{\left(\omega_{k}\mathopen{}\mathclose{{\left(t-\tfrac{1}{2}}}\right)}}\right)+\sum_{j=1}^{N_{\mathrm{basis}}}\frac{\sigma_{k}}{\bigl((j\pi)^{2}+\tau_{k}^{2}\bigr)^{\alpha_{k}/2}}z_{j}\sqrt{2}\sin(j\pi t)\,.$ (F.12)
@@
We view $r\colon[0,1]\to\partial\Omega$ as a 1D parametrization of the boundary $\partial\Omega\subset\mathbb{R}^{2}$ for the domain $\Omega\coloneqq(0,1)^{2}$. The entire domain is discretized on an $M\times M$ grid, where $M$ is as in (F.11). Here $z_{j}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$, $\alpha_{k}=1.5$ for all $k$, and $N_{\mathrm{basis}}=4M-4$. We set the grid parameter to be $M=128$ and sample $K=64$ test distributions in this example. The parameters $\omega_{k}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\rho([-2\pi,2\pi])$ and $\tau_{k}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\rho([0,50])$ are sampled from a custom distribution $\rho=\rho([x_{\min},x_{\max}])$ and held fixed throughout the algorithm. The custom distribution $\rho$ has the probability density function
$p(x)=C\sin\bigl(x-\tfrac{\pi}{2}\bigr)+1$ (F.13)
@@
for $x\in[x_{\min},x_{\max}]\subset\mathbb{R}$, where $C$ is a normalizing constant ensuring that $p$ integrates to $1$ over the domain $x_{\min},x_{\max}]$. Samples are generated by using inverse transform sampling. Any time a Neumann boundary condition $g$ is sampled according to (F.12) and (F.13), we subtract the spatial mean of $g$ from itself to ensure that the resulting function integrates to zero. This is necessary to make the Neumann PDE problem—as well as the NtD operator itself—well-posed.
@@
$t\mapsto\widetilde{g}(t)=10\sin\mathopen{}\mathclose{{\left(\omega\mathopen{}\mathclose{{\left(t-\tfrac{1}{2}}}\right)}}\right)+\sum_{j=1}^{N_{\mathrm{basis}}}\frac{\sigma}{\bigl((j\pi)^{2}+\tau^{2}\bigr)^{\alpha/2}}\widetilde{z}_{j}\sqrt{2}\sin(j\pi t)\,$, (F.14)
@@
Note that $g$ and $\widetilde{g}$ are finite-term truncations of the KLE of Gaussian measures with Matérn-like covariance operators; recall Lem. B.2. In particular, for $\widetilde{g}$, and similarly for $g$, the covariance operator of the Gaussian measure has eigenpairs
$\lambda_{j}=\frac{\sigma^{2}}{\bigl((j\pi)^{2}+\tau^{2}\bigr)^{\alpha}}\quad\text{and}\quad t\mapsto\phi_{j}(t)=\sqrt{2}\,\sin(j\pi t)\quad\text{for}\quad j=1,2,\dots,N_{\mathrm{basis}}\,.$ (F.15)
Using (F.15) and (B.1), the squared $2$-Wasserstein distance in Prop. 3.1 reduces to
$\mathsf{W}_{2}^{2}(\nu,\nu_{k}^{\prime})=\int_{0}^{1}\Bigl(10\sin\mathopen{}\mathclose{{\left(\omega\mathopen{}\mathclose{{\left(t-\tfrac{1}{2}}}\right)}}\right)-10\sin\mathopen{}\mathclose{{\left(\omega_{k}\mathopen{}\mathclose{{\left(t-\tfrac{1}{2}}}\right)}}\right)\Bigr)^{2}dt+\sum_{j=1}^{N_{\mathrm{basis}}}\Bigl(\sqrt{\lambda_{j}}-\sqrt{\lambda_{j,k}}\Bigr)^{2},$ (F.16)
@@
$\mathsf{m}_{2}(\nu)=\int_{0}^{1}\Bigl(10\sin\mathopen{}\mathclose{{\left(\omega\mathopen{}\mathclose{{\left(t-\tfrac{1}{2}}}\right)}}\right)\Bigr)^{2}dt+\sum_{j=1}^{N_{\mathrm{basis}}}\lambda_{j}\,.$ (F.17)
These closed-form reductions follow directly from the finite-dimensional KLE and Ex. B.1.
@@
$\max_{(g_{1},g_{2})\in\mathsf{S}}\frac{\lVert\mathcal{G}(g_{1})-\mathcal{G}(g_{2})\rVert_{L^{2}(\partial\Omega)}}{\lVert g_{1}-g_{2}\rVert_{L^{2}(\partial\Omega)}}$ (F.18)
@@
The model architecture is a DeepONet comprising a branch network with five fully connected layers of sizes (4M4)(4M-4), $512$, $512$, $512$, and $200$, and a trunk network with five fully connected layers of sizes $2$, $128$, $128$, $128$, and $200$, respectively; see SM E.1 for details. All layers use ReLU activation functions. Weights are initialized using the Glorot normal initializer. The model is trained with the Adam optimizer with a learning rate of $10^{-3}$. After the model-training half step in the alternating minimization scheme Alg. 2, the parameter $\omega$ in (F.14) is optimized using SciPy’s default implementation of differential evolution with a tolerance of $10^{-7}$. Due to the presence of many local minima in the optimization landscape, this global optimization algorithm is necessary.
@@
The initial value for $\omega$ is set to $8\pi$. For every model-training step, $500$ pairs of Neumann and Dirichlet data are generated using the current value of $\omega$. For practical reasons, the model is initially trained for $2000$ iterations of Adam. If the AMA loss has not improved relative to the previous iteration, training continues for an additional $2000$ iterations. This process is repeated up to $10$ times, after which the algorithm proceeds to the parameter optimization step for the training distribution $\nu$. During distribution parameter optimization, the AMA loss is evaluated using $500$ newly generated data pairs, used specifically to compute the first term of the objective (4.7). To ensure deterministic behavior, the $500$ sets of random coefficients $\{z_{j}\}_{j=1}^{N_{\mathrm{basis}}}$ are fixed separately for training and optimization. This results in $1000$ total sets of coefficients fixed for the entire algorithm. This process is repeated across $80$ independent runs using the same conductivity but different realizations of the random coefficients $\{z_{j}\}$. The relative AMA loss and relative OOD error is computed for each run; see (F.19).
@@
$\text{Relative OOD Error}\coloneqq\sqrt{\frac{\operatorname{\mathbb{E}}_{\nu^{\prime}\sim\mathbb{Q}}\operatorname{\mathbb{E}}_{g\sim\nu^{\prime}}\|\mathcal{G}^{\star}g-\mathcal{G}_{\theta}(g)\|_{L^{2}}^{2}}{\operatorname{\mathbb{E}}_{\nu^{\prime}\sim\mathbb{Q}}\operatorname{\mathbb{E}}_{g\sim\nu^{\prime}}\|\mathcal{G}^{\star}g\|_{L^{2}}^{2}}}\,$, (F.19)
@@
$\text{Relative OOD Error}\approx\sqrt{\frac{\frac{1}{K}\sum_{k=1}^{K}\frac{1}{N}\sum_{n=1}^{N^{\prime}}\|\mathcal{G}^{\star}g_{k,n}-\mathcal{G}_{\theta}(g_{k,n})\|_{L^{2}}^{2}}{\frac{1}{K}\sum_{k=1}^{K}\frac{1}{N^{\prime}}\sum_{n=1}^{N^{\prime}}\|\mathcal{G}^{\star}g_{k,n}\|_{L^{2}}^{2}}}\,$, (F.20)
@@
The computational resources here are the same as those reported in SM F.2, except three total CPU and GPU hours were needed for all independent runs.
F.4 Darcy flow forward operator learning
@@
In this example from Sec. 5, we are interested in learning another operator that is relevant and related to the EIT problem. In contrast with the previous example in SM F.3 where we fix the conductivity, the Darcy flow forward operator that we consider is nonlinear and maps the log-conductivity function $\log(a)$ to the PDE solution $u$ of
$\mathopen{}\mathclose{{\left\{\begin{aligned} -\nabla\cdot(a\nabla u)&=1&&\quad\text{in}\quad\Omega\,,\\ u&=0&&\quad\text{on}\quad\partial\Omega\,.\end{aligned}}}\right$. (F.21)
That is, $\mathcal{G}^{\star}\colon\log(a)\mapsto u$. Here, $\Omega\coloneqq(0,1)^{2}$.
Again, we first assign a random measure $\mathbb{Q}$ to evaluate the OOD performance. We take $\mathbb{Q}=\frac{1}{K}\sum_{k=1}^{K}\delta_{\nu_{k}^{\prime}}$, where each test distribution $\nu_{k}^{\prime}$ is defined by parameters sampled i.i.d. from custom distributions (F.13): $m_{k}\sim\rho([0,1])$, $\alpha_{k}\sim\rho([2,3])$, and $\tau_{k}\sim\rho([1,4])$. The scaling parameter is set as $\sigma_{k}=\tau_{k}^{\alpha_{k}-1}$. We set $K=20$ and the grid parameter $M=64$. Given these parameters, each random log-conductivity function $\log(a^{\prime})$ from test distribution $\nu_{k}^{\prime}$ has the form
$x\mapsto\log(a^{\prime}(x))=m_{k}+\sum_{i=1}^{M+1}\sum_{j=1}^{M+1}\frac{2\sigma_{k}}{\bigl((i\pi)^{2}+(j\pi)^{2}+\tau_{k}^{2}\bigr)^{\alpha_{k}/2}}z^{\prime}_{ij}\sin(i\pi x_{1})\sin(j\pi x_{2})\,$, (F.22)
where $z^{\prime}_{ij}\sim\mathcal{N}(0,1)$ independently. Thus, $a$ is lognormal. Sample log-conductivities $\log(a)$ from training distribution $\nu$ are assumed to follow the same form
$x\mapsto\log(a(x))=m+\sum_{i=1}^{M+1}\sum_{j=1}^{M+1}\frac{2\sigma}{\bigl((i\pi)^{2}+(j\pi)^{2}+\tau^{2}\bigr)^{\alpha/2}}z_{ij}\sin(i\pi x_{1})\sin(j\pi x_{2})\,$, (F.23)
with fixed parameters $\alpha=2$, $\tau=3$, and $\sigma=3$ and $z_{ij}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. Alg. 2 is used to optimize the mean shift parameter $m\in\mathbb{R}$ in order to reduce the OOD error. As in SM F.3, $\log(a)$ is modeled via a finite KLE of a Gaussian measure with Matérn-like covariance operator. Thus, by working with $\log(a)$ directly, we can exploit (B.1) to write the squared $2$-Wasserstein distance factor in objective (4.7) in closed-form.
Model architecture and optimization algorithm.
The same DeepONet architecture and training setup from SM F.3 are used here, except the branch network has layer sizes $M^{2}$, $256$, $256$, $128$, and $200$. After the model-training half step in Alg. 2, the scalar parameter $m$ is optimized using PyTorch’s Adam optimizer with a learning rate of $10^{-3}$.
Training and optimization procedure.
The initial value of $m$ is set to zero. The training and optimization procedure follows the same structure as in SM F.3, but with the following differences: $200$ conductivity-solution pairs are generated for training using the current value of $m$. The model is initially trained for $5000$ iterations. If the AMA loss has not improved relative to the previous iteration, training continues for an additional $5000$ iterations. This process is repeated up to three times, after which the algorithm proceeds to the distribution parameter optimization step. During distribution parameter optimization, $500$ steps are taken. Then the AMA loss is evaluated using three newly generated data pairs. If the AMA loss has not improved relative to the previous iteration, another $500$ steps are taken, up to $1500$ total steps. After this, the procedure returns to model training again. To ensure deterministic behavior, $203$ sets of random coefficients $\{z_{ij}\}$ are fixed across training and optimization. The preceding process is repeated across $13$ independent runs, each using different realizations of the coefficients. The average relative OOD error and average relative AMA loss are computed for each run.
Model evaluation.
After $13$ independent runs, the average relative OOD error and relative AMA loss are computed and are shown in Fig. 16. In Fig. 4, we visualize the pointwise absolute error for iterations $1$, 22, 33, and $8$. At these iterations, the DeepONet model is trained using the optimal training distribution produced by AMA. The model is then evaluated on the same fixed test conductivity. The progressive reduction in the spatial error across these panels of Fig. 4 corroborates the quantitative decrease in the relative OOD error, which converges to approximately $4\%$ in our final model.
Figure 16: Average relative OOD error and average relative loss of Alg. 2 on the Darcy flow forward problem. (Left) After $13$ independent runs, decay of the average relative OOD error of the model when trained on the optimal distribution identified at each iteration; a 9595% confidence interval of the true relative OOD error is provided at each iteration. (Right) Decay of average AMA loss defined in (4.7) vs. iteration relative to the same loss at initialization.
Computational resources.
The computational resources here are the same as those reported in SM F.3, except three total CPU and GPU hours were needed for all independent runs.
F.5 Radiative transport operator learning
In this example from Sec. 5, we consider the hyperbolic PDE for radiative transport governing the particle density $u(z,v)$ at position $z\in\Omega\subset\mathbb{R}^{d}$ and velocity $v\in V\subset\mathbb{R}^{d}$: